# bibliography.bib

### v1

@@ Abstract
This study presents findings from long-term biometric evaluations conducted at the Biometric Evaluation Center (BEZ). Over the course of two and a half years, our ongoing research with over 400 participants representing diverse ethnicities, genders, and age groups were regularly assessed using a variety of biometric tools and techniques at the controlled testing facilities. Our findings are based on the General Data Protection Regulation (GDPR)- compliant local BEZ database with more than 238.000 biometric data sets categorized into multiple biometric modalities such as face and finger. We used state-of-the-art face recognition (FR) algorithms to analyze long-term comparison scores. Our results show that these scores fluctuate more significantly between individual days than over the entire measurement period. These findings highlight the importance of testing biometric characteristics of the same individuals over a longer period of time in a controlled measurement environment and lays the groundwork for future advancements in biometric data analysis.
@@ 1 Introduction
A long-term evaluation approach offers a solution to this problem by enabling systematic and controlled testing over extended periods. By maintaining controlled and adjustable conditions, several key advantages can be realized. First, the same systems can be tested with the same individuals over time, enabling robust longitudinal analysis, to evaluate the aging processes of people and systems as well as seasonal dependencies. Second, different systems can be evaluated either simultaneously or within a short time frame under identical conditions, which facilitates direct comparisons. Third, newer versions of a system can be assessed in relation to their predecessors, providing insights into technological progress. In addition, extensions and modifications of existing systems, as well as the development of entirely new systems, can be evaluated in a structured manner. Finally, usability aspects can be examined thoroughly under these consistent conditions.
@@ 3.1 BEZ evaluation process
The biometric database used in this study was constructed based on the collected data from recruited participants under controlled conditions. The GDPR compliant evaluation process is divided into two main phases, as illustrated in Figure 1: First, each subject underwent an initial enrollment phase to establish reference biometric data and to familiarize themselves with the study procedures. At the beginning of the enrollment session, participants attended a brief introductory session, which included a short guided tour of the testing environment and a detailed explanation of the study’s objectives, methodology and privacy policies. Then participants signed an informed consent form, ensuring voluntary participation and compliance with ethical research guidelines. Within this process, every proband is issued with an unique ID to pseudonymize identities. Afterwards, access to the connection between ID and real name is only accessible for the Federal Office for Information Security (BSI) ’s data protection officer, who also administers the declarations of consents. Thus, the BEZ databases only store the pseudonymized IDs along with the recorded data. The test subjects can at any time request information about their stored data as well as the deletion of this data, including any backups, from the data protection officer. Timely processing will then be ensured.
During the enrollment a high resolution and high quality set of facial images is captured with the K13 photo studio to serve as the initial reference for FR. The K13 photo studio offers standardized lighting, uniform background and positional conditions. Each record includes high quality images from 13 cameras (see Figure 2). With camera 5, a full frontal image of the face is captured. Eight additional cameras are positioned in angles of 15 degree, whereas the cameras on the diagonal have angles of 15° both on horizontal and vertical axes. The outer four cameras take images in 45° positions.
@@
Additionally all fingers of the left and right hand are captured during the enrollment using a high-resolution optical four-finger fingerprint scanner. Each finger is captured in single capture mode. These recorded reference facial image and fingerprints as well as the proband data (age, gender, etc.) and consents given are entered into the reference database. The stored reference data records of the test subjects are updated in a regular manner.
After enrollment, in the second phase the main data acquisition process is carried out. Probands can use the self-service kiosks for autonomous check in to the BEZ evaluations. The kiosk first captures a facial image, which is then compared against the reference database of enrolled participants. In case of a single match, the subject is identified and the unique proband ID is found. The kiosk writes this ID together with the facial reference image from the database to a smartcard. This smartcard can then be used at the biometric data acquisition stations where proband ID is connected to all recordings of this person. Furthermore, due to the stored reference image, the smartcard can be used like a passport to enter the experimental automated border control (ABC) gates. The recorded biometric data together with experiment specific meta-information (e.g. subject ID, experiment, time, system-settings, environmental conditions, specific headgear and glasses…) is saved to the BEZ database. To ensure easy handling and high-quality dataset curation, biometric samples of each experiment station is annotated by a custom labeling scheme. After evaluation, probands can go back to the self-services kiosk to check out. The subject receives a credit note for the participation and can have vouchers issued after several visits.
@@ 3.2 Datasets
With the above acquisition approach we recruited 492 participants, which have finished the enrollment process. The participant pool consists of 254 male, 180 female, 1 diverse and 57 not specified individuals (see Figure 3). 428 have specified their birthdate or year of birth. The average age of these probands is 30.6 ± 11.4. The oldest individual is 89 years old, the lowest specified age is 4. The individual with the age of 4 is evidently a wrong statement in the signed consent form of that person. With the corrected dataset the youngest individual is 18 years old. Two pairs of biological identical twins participated in the evaluations.
@@
The evaluations took place three days a week in time slots of 2.5 hours. On average, around 96 participants took part in the evaluations per week and around 32 per evaluation day, whereby these figures can vary greatly depending on the time of year, among other things due to vacation time, lecture-free period.
Each participant provided data across several sessions on multiple experiment stations to account for temporal variations. These stations were arranged to facilitate smooth participant flow and allow for simultaneous recordings and minimizing waiting times. The collected datasets (see Table 1) include multiple biometric modalities, each with specific capture devices and recording conditions. At the already mentioned K13 photo studio, high-quality images from 13 perspectives of the subjects were recorded. In addition, ABC gates operated as at major national airports as well as an experimental gate and a stand-alone gate panel are in use. Long-term test series are of particular importance for these ABC-gates, as the effects of changing characteristics (skin, hairstyle, beard) pose a challenge, especially in view of the long validity of passports. At the fingerprint stations we alternated between optical and capacitive fingerprint scanners and fingerprints have been captured from both hands in flat and rolled mode. Additionally skin reflectance and transmission values have been captured at the visible and near-infrared (NIR) light spectrometer and the unique hardware presentation attack detection (PAD) system LokiMk2 [scheer_customizable_2024] station. At the thermal camera station, different camera systems captured thermal images of the subjects face. Another camera station used the Time-of-Flight (ToF) technology to capture 3D face data of subjects and is currently combined with remote photoplethysmography (rPPG). In addition to the long term tests, there were also individual sessions, so-called mass tests, in which special stations were set up and tested. These include the optical coherence tomography (OCT) experiment [kirfel2022robust] where 3D fingerprints have been captured and an experiment with the focus on the back of the hand, where a camera and light system has been setup, to capture high-quality images of the back of the hand in different hand gestures. To evaluate the usability and human-technology interaction under realistic conditions, a dedicated study was conducted at the ABC-gates as part of a research project at the BEZ [paul2022usability].
@@ 3.3 Analysis
The focus of this study is limited to the long-term effects of face images in regards to recognition accuracy, which is examined across different FR algorithms. Face comparison scores have been obtained from COTS1, COTS2, ArcFace [deng_arcface_2018], AdaFace [kim_adaface_2022] and DLIB FR [e_king_dlib-ml_2009, parkhi_deep_2015] to allow for comparison in-between different algorithms.
@@ 3.4 Definitions
The following terminology corresponding to the ISO 30107 [international_organization_for_standardization_information_2023], ISO/IEC 2382-37 [international_organization_for_standardization_information_2022] and the BSI TR-03166 [bsi_bsi_2024] is defined for this study. The term “biometric score” refers to an evaluation of the comparison of two facial images by a FR algorithm and is therefore a way of estimating the similarity of two faces. If the score exceeds a threshold value set for the respective application, it is assumed that both images originate from the same person (“mated”). In the FR algorithms used, the biometric score is specified in an interval between ”0” and ”1”, with ”1” corresponding to the best possible match between the two facial images.
@@
Additionally, the False Match Rate (FMR) is defined as the percentage of biometric comparisons of different identities that incorrectly result in a match. The False Non-Match Rate (FNMR) is accordingly the proportion of mated comparisons with a biometric score above the recognition threshold.
@@ 4.2 Biometric Performance
As a first evaluation of the recorded data and a benchmark for the longitudinal studies we investigated the performance of commercial off-the-shelf (COTS) and Open-Source [deng_arcface_2018, e_king_dlib-ml_2009, parkhi_deep_2015, kim_adaface_2022] FR algorithms on the K13 dataset. Initially, using only the frontal images of camera 5 for the D_3.1 dataset (see Table 2) results in the biometric score distribution shown in Figure 4 and the corresponding statistics summarized in Tables 3 and 4. As expected of a high quality data set in terms of the record image quality and the data labeling, a sharp separation between mated and non-mated comparison scores can be found. The EER of $3.94\cdot 10^{-6}$  is achieved for a threshold of $0.92$ here (see Figure 5). However, for a real application scenario the threshold would be chosen quite lower to reduce false rejections due to potentially variations and lower qualities than these highly controlled, high quality K13 images. A threshold of $T=0.755$ seems to be more realistic as it is the lower boundary that would lead to the same error rates and the same number of outliers: still no false rejections and only $24$ false acceptances occur. A closer look to these $24$ outlier comparisons shows, that the compared subjects are biological identical twins, which look very similar even for a human operator. The mean mated score for these twins is $0.9938\pm 0.0047$, whereas the mean non-mated score is $0.9640\pm 0.0067$. Therefore, differentiability based on this comparison algorithm seems theoretically possible in this special case with the high quality K13 data with minimal variations.
@@ 4.3 Angle Analysis
In the next study, the full frontal K13 images of camera 5 (see Figure 2) was compared to the additional cameras, whereby only comparisons to the different angles of the same recordings were considered. By this method only the influence of the angle is decisive and not the variance of different images of the same person. We define the x-angle as the rotation around the subject on a horizontal plane (left-right movement) and y-angles as the vertical elevation respectively (up-down movement). Figures 6 and 7 as well as Tables 6 and 7 show the results for various x- (horizontal) and y-angles (vertical), whereas the respective other angle is 0. Thus, influence of horizontal and vertical angles on the biometric scores can be investigated separately. Angle 0 would accordingly be a comparison to the same image and is therefore not plotted here.
@@ 4.4 Longterm Analysis
To analyze the temporal stability of facial recognition performance, delta scores were plotted against the number of days since the initial image acquisition for all participants using the COTS and open-source algorithms. As shown in Figure 8, each point represents the delta score for an individual genuine comparison relative to the first capture.
@@
While individual variations exist and intra-day similarity scores fluctuate stronger than inter-day similarity scores, the general trend suggests a consistent and measurable temporal drift in facial features. The distribution is densest in the early time intervals, as can be seen in figure 9. Here delta scores tend to cluster closer to zero, whereas longer intervals are associated with more negative delta scores. This aligns with expectations from facial biometric systems, where changes in appearance due to aging or other temporal factors can reduce match confidence over time.
A similar analysis was conducted for the remaining four facial recognition algorithms as seen in figure 10. Across all algorithms, a general decline in delta scores over time was observed. However, the magnitude of this decline varied significantly between algorithms. The effect on the delta score is not directly comparable between the algorithms, as their functionality differs from each other. Therefore for each algorithm a separate threshold has been calculated. The open-source algorithms exhibit a noticeably stronger negative trend in their delta scores compared to the COTS algorithms. This suggests that these algorithms may be more sensitive to temporal variation and possibly less robust in handling aging-related facial changes or other longitudinal appearance shifts.
@@ 5 Discussion
The dataset acquired through this long-term study offers a unique opportunity for biometric research, particularly in assessing the stability and variance of facial biometric data over time. Our findings reveal that temporal changes in facial biometric scores over 2.5 years are minimal in COTS algorithms, with inter-individual score distances remaining significantly larger than intra-individual temporal variations. This underscores the reliability of facial biometrics as a stable identifier over time.

### v2

@@ Abstract
This study presents findings from long-term biometric evaluations conducted at the Biometric Evaluation Center (BEZ). Over the course of more than two and a half years, our ongoing research with over 400 participants representing diverse ethnicities, genders, and age groups were regularly assessed using a variety of biometric tools and techniques at the controlled testing facilities. Our findings are based on the General Data Protection Regulation (GDPR)- compliant local BEZ database with more than 238,000 biometric data sets categorized into multiple biometric modalities such as face and finger. We used state-of-the-art face recognition (FR) algorithms to analyze long-term comparison scores. Our results show that these scores fluctuate more significantly between individual days than over the entire measurement period. These findings highlight the importance of testing biometric characteristics of the same individuals over a longer period of time in a controlled measurement environment and lays the groundwork for future advancements in biometric data analysis.
@@ 1 Introduction
A long-term evaluation approach offers a solution to this problem by enabling systematic and controlled testing over extended periods. By maintaining controlled and adjustable conditions, several key advantages can be realized. First, the same systems can be tested with the same individuals over time, enabling robust longitudinal analysis, to evaluate the aging processes of people and systems. Second, different systems can be evaluated either simultaneously or within a short time frame under identical conditions, which facilitates direct comparisons. Third, newer versions of a system can be assessed in relation to their predecessors, providing insights into technological progress. In addition, extensions and modifications of existing systems, as well as the development of entirely new systems, can be evaluated in a structured manner. Finally, usability aspects can be examined thoroughly under these consistent conditions.
@@ 3.1 Participant acquisition
To create an additional incentive for participation in the evaluations, vouchers are issued to the subjects after several visits.
3.2 BEZ evaluation process
@@
The biometric database used in this study was constructed based on the collected data from recruited participants under controlled conditions. The GDPR compliant evaluation process is divided into two main phases, as illustrated in Fig. 1: First, each subject underwent an initial enrollment phase to establish reference biometric data and to familiarize themselves with the study procedures. At the beginning of the enrollment session, participants attended a brief introductory session, which included a short guided tour of the testing environment and a detailed explanation of the study’s objectives, methodology and privacy policies. Then participants signed an informed consent form, ensuring voluntary participation and compliance with ethical research guidelines. Within this process, every proband is issued with an unique ID to pseudonymize identities. Afterwards, access to the connection between ID and real name is only accessible for the Federal Office for Information Security (BSI) ’s data protection officer, who also administers the declarations of consents. Thus, the BEZ databases only store the pseudonymized IDs along with the recorded data. The test subjects can at any time request information about their stored data as well as the deletion of this data, including any backups, from the data protection officer. Timely processing will then be ensured.
During the enrollment a high resolution and high quality set of facial images is captured with the K13 photo studio to serve as the initial reference for FR. The K13 photo studio offers standardized lighting, uniform background and positional conditions. Each record includes high quality images from 13 cameras (see Fig. 2). With camera 5, a full frontal image of the face is captured. Eight additional cameras are positioned in angles of 15 degree, whereas the cameras on the diagonal have angles of 15° both on horizontal and vertical axes. The outer four cameras take images in 45° positions.
@@
Additionally all fingers of the left and right hand are captured during the enrollment using a high-resolution optical four-finger fingerprint scanner. Each finger is captured in single finger capture mode. These recorded reference facial image and fingerprints as well as the proband data (age, gender, etc.) and consents given are entered into the reference database. The stored reference data records of the test subjects are updated in a regular manner.
After enrollment, in the second phase the main data acquisition process is carried out. Probands can use the self-service kiosks for autonomous check in to the BEZ evaluations. The kiosk first captures a facial image, which is then compared against the reference database of enrolled participants. In case of a single match, the subject is identified and the unique proband ID is found. The kiosk writes this ID together with the facial reference image from the database to a smartcard. This smartcard can then be used at the biometric data acquisition stations where proband ID is connected to all recordings of this person. Furthermore, due to the stored reference image, the smartcard can be used like a passport to enter the experimental automated border control (ABC) gates. The recorded biometric data together with experiment specific meta-information (e.g. subject ID, experiment, time, system-settings, environmental conditions, specific headgear and glasses…) is saved to the BEZ database. To ensure easy handling and high-quality dataset curation, biometric samples of each experiment station is annotated by a custom labeling scheme. After evaluation, probands can go back to the self-services kiosk to check out, whereby all data on the used smartcard is automatically deleted. The subject receives a credit note for the participation.
@@ 3.3 Datasets
With the above acquisition approach we recruited 492 participants, which have finished the enrollment process. The participant pool consists of 254 male, 180 female, 1 diverse and 57 not specified individuals (see Fig. 3). 428 have specified their birth date or year of birth. The average age of these probands is 30.6 ± 11.4. The oldest individual is 89 years old, the lowest specified age is 18. Two pairs of biological identical twins participated in the evaluations.
@@
The evaluations took place three days a week in time slots of 2.5 hours. On average, around 96 participants took part in the evaluations per week and around 32 per evaluation day, whereby these figures can vary greatly with up to 146 visits per week depending on the time of year, among other things due to vacation time, lecture-free period.
Each participant provided data across several sessions on multiple experiment stations. These stations were arranged to facilitate smooth participant flow and allow for simultaneous recordings and minimizing waiting times. The collected datasets (see Table 1) include multiple biometric modalities, each with specific capture devices and recording conditions. At the already mentioned K13 photo studio, high-quality images from 13 perspectives of the subjects were recorded. In addition, ABC gates operated as at major national airports as well as an experimental gate and a stand-alone gate panel are in use. Long-term test series are of particular importance for these ABC-gates, as the effects of changing characteristics (skin, hairstyle, beard) pose a challenge, especially in view of the long validity of passports. At the fingerprint stations we alternated between optical and thin film transistor (TFT) based fingerprint scanners and fingerprints have been captured from both hands in flat and rolled mode. Additionally the remission of skin is evaluated by capture of spectra in the visible and near-infrared (NIR) range and the unique hardware presentation attack detection (PAD) system LokiMk2 [scheer_customizable_2024] station. At the thermal camera station, different camera systems captured thermal images of the subjects face. Another camera station used the Time-of-Flight (ToF) technology to capture 3D face data of subjects and is currently combined with remote photoplethysmography (rPPG). In addition to the long term tests, there were also individual sessions, so-called mass tests, in which special stations were set up and tested. These include the optical coherence tomography (OCT) experiment [kirfel2022robust] where 3D fingerprints have been captured and an experiment with the focus on the back of the hand, where a camera and light system has been setup, to capture high-quality images of the back of the hand in different hand gestures. To evaluate the usability and human-technology interaction under realistic conditions, a dedicated study was conducted at the ABC-gates as part of a research project at the BEZ [paul2022usability].
@@ 3.4 Analysis
The focus of this study is limited to the long-term effects of face images in regards to recognition accuracy, which is examined across different FR algorithms. Face comparison scores have been obtained from commercial off-the-shelf (COTS) algorithmns, here called COTS1 and COTS2, as well as for the open-source algorithms ArcFace [deng_arcface_2018], AdaFace [kim_adaface_2022] and DLIB FR [e_king_dlib-ml_2009, parkhi_deep_2015] to allow for comparison in-between different software.
@@ 3.5 Definitions
The following terminology corresponding to the ISO 30107 [international_organization_for_standardization_information_2023], ISO/IEC 2382-37 [international_organization_for_standardization_information_2022] and the BSI TR-03166 [bsi_bsi_2024] is defined for this study. The term “comparison score” refers to an evaluation of the comparison of two facial images by a FR algorithm and is therefore a way of estimating the similarity of two faces. If the score exceeds a threshold value set for the respective application, it is assumed that both images originate from the same person (“mated”). In the FR algorithms used, the comparison score is specified in an interval between ”0” and ”1”, with ”1” corresponding to the best possible match between the two facial images.
@@
Additionally, the False Match Rate (FMR) is defined as the percentage of biometric comparisons of different identities that incorrectly result in a match. The False Non-Match Rate (FNMR) is accordingly the proportion of mated comparisons with a comparison score above the recognition threshold.
@@ 4.2 Biometric Performance
As a first evaluation of the recorded data and a benchmark for the longitudinal studies we investigated the performance of COTS and open-source [deng_arcface_2018, e_king_dlib-ml_2009, parkhi_deep_2015, kim_adaface_2022] FR algorithms on the K13 dataset. Initially, using only the frontal images of camera 5 for the D_3.1 dataset (see Table 2) results in the comparison score distribution shown in Fig. 4 and the corresponding statistics summarized in Tables 3 and 4. As expected of a high quality data set in terms of the record image quality and the data labeling, a sharp separation between mated and non-mated comparison scores can be found. The EER of $3.94\cdot 10^{-6}$  is achieved for a threshold of $0.92$ here (see Fig. 5). However, for a real application scenario the threshold would be chosen quite lower to reduce false rejections due to potentially variations and lower qualities than these highly controlled, high quality K13 images. A threshold of $T=0.755$ seems to be more realistic as it is the lower boundary that would lead to the same error rates and the same number of outliers: still no false rejections and only $24$ false acceptances occur. A closer look to these $24$ outlier comparisons shows, that the compared subjects are biological identical twins, which look very similar even for a human operator. The mean mated score for these twins is $0.9938\pm 0.0047$, whereas the mean non-mated score is $0.9640\pm 0.0067$. Therefore, differentiability based on this comparison algorithm seems theoretically possible in this special case with the high quality K13 data with minimal variations.
@@ 4.3 Angle Analysis
In the next study, the full frontal K13 images of camera 5 (see Fig. 2) was compared to the additional cameras, whereby only comparisons to the different angles of the same recordings were considered. By this method only the influence of the angle is decisive and not the variance of different images of the same person. We define the x-angle as the rotation around the subject on a horizontal plane (left-right movement) and y-angles as the vertical elevation respectively (up-down movement). Figures 6 and 7 as well as Tables 6 and 7 show the results for various x- (horizontal) and y-angles (vertical), whereas the respective other angle is 0. Thus, influence of horizontal and vertical angles on the comparison scores can be investigated separately. Angle 0 would accordingly be a comparison to the same image and is therefore not plotted here.
@@ 4.4 Longterm Analysis
To analyze the temporal stability of facial recognition performance, delta scores were plotted against the number of days since the initial image acquisition for all participants using the COTS and open-source algorithms. As shown in Fig. 8, each point represents the delta score for an individual genuine comparison relative to the first capture.
@@
While individual variations exist and intra-day similarity scores fluctuate stronger than inter-day similarity scores, the general trend suggests a consistent and measurable temporal drift in facial features. The distribution is densest in the early time intervals, as can be seen in Fig. 9. Here delta scores tend to cluster closer to zero, whereas longer intervals are associated with more negative delta scores. This aligns with expectations from facial biometric systems, where changes in appearance due to aging or other temporal factors can reduce match confidence over time.
A similar analysis was conducted for the remaining four facial recognition algorithms as seen in Fig. 10. Across all algorithms, a general decline in delta scores over time was observed. However, the magnitude of this decline varied significantly between algorithms. The effect on the delta score is not directly comparable between the algorithms, as their functionality differs from each other. Therefore for each algorithm a separate threshold has been calculated. The open-source algorithms exhibit a noticeably stronger negative trend in their delta scores compared to the COTS algorithms. This suggests that these algorithms may be more sensitive to temporal variation and possibly less robust in handling aging-related facial changes or other longitudinal appearance shifts.
@@ 5 Discussion
The dataset acquired through this long-term study offers a unique opportunity for biometric research, particularly in assessing the stability and variance of facial biometric data over time. Our findings reveal that temporal changes in facial comparison scores over 2.5 years are minimal in COTS algorithms, with inter-individual score distances remaining significantly larger than intra-individual temporal variations. This underscores the reliability of facial biometrics as a stable identifier over time.